A company issues P20,000,000, 7.8%, 20-year bonds to yield 8% on January 1, 2019. Interest is paid on June 30 and December 31. The proceeds from the bonds are P19,604,145. Using effective-interest amortization, how much interest expense will be recognized in 2019?
Answer #1

Interest expense on June 30 = issued price * Market rate of interest * 6 / 12

Interest expense on June 30 = P19,604,145 * 8% * 6 / 12 = P784,165.80

Carrying value of bonds at the end of June 30 = P19,604,145 - [(P20,000,000*7.8% * 6/12) - P784,165.80]

Carrying value of bonds at the end of June 30 = P19,608,310.80

Interest expense on December 31 = P19,608,310.80 * 8%*6/12 = P784,332.432

Total interest expense to be recognized in 2019 is P784,165.80 + P784,332.432 = P1,568,498.232 or 1,568,498 (rounded off to nearest whole number)